# Numerical model for pellet rocket acceleration in PELOTON

### v1

@@ 4.3.1 Plasma gradients
Using PELOTON’s plasma cooling module, we computed plasma states after the injection of first 30 fragments (see Figure 5(b)), about 10 percent of a typical ITER injection, and used these states to estimate how plasma gradients affect the rocket acceleration. As we observe, after ablated material is deposited the plasma electron temperature remains practically constant, approximately 600 eV, within normalized flux coordinates from 0.9 to 0.5; plasma density also does not experience large gradients. As a result, plasma gradients changes the rocket acceleration by negligibly small amounts in this domain. Therefore, we estimate the upper bound of the plasma gradient effect by considering largest plasma gradients at the edge of the ablated material deposition with the flux coordinate of about 0.5.
(b) (a) Plasma electron density, $m^{-3}$ and (b) temperature (eV) after the injection of 30 SPI fragments.
@@ 5 Summary and future work
[…]
This work has been supported in part by the ITER Organization and the ITER Disruption Mitigation Task Force via contract. The views and opinions expressed herein do not necessarily reflect those of the ITER Organization. The Swiss contribution was supported in part by the Swiss National Science Foundation.

### v2

@@ 4.3 Influence of neon component
4.4 Influence of Magnetic Field Strength and Radius of Curvature
In this section we report results from a parameter scan about one simulation point at JET conditions. In particular, we select point 20 from table 4 in the appendix and perform four simulations, each of which varies one parameter at a time. We vary the magnetic field strength B and the pellet’s radial coordinate R. Simulation point 20 used B = 2.8 T and R = 3 m for a 1 mm pellet with background plasma with electron temperature and density at approximately 1220 eV and 1.25e20 $m^{-3}$, respectively. As such, varying one parameter at a time gives simulations at B = 3.8 T, B = 1.8 T, R = 4.0 m, and R = 2.0 m, where we have selected to vary B by 1 T and R by 1 m. The results are presented in tables 1 and 2. Varying the R coordinate did not have a large impact on the ablation rate or the pressure asymmetry leading to the rocket acceleration. We observed that the ablation rate slightly decreased with increases in the magnetic field, and that the pressure asymmetry increased with increasing magnetic field strength.
Radial Coordinate (m) G (g/s) $\Delta P$ (bar)
4.0 10.2 0.70
3.0 10.3 0.68
2.0 10.4 0.6
Table 1: We perform simulations in which we vary the radial coordinate by 1 m, leaving all other simulation parameters the same as in point 20 in table 4. Point 20 is represented by the middle row. In particular, these simulations use a 1 mm pellet with background plasma at approximately electron temperature 1220 eV and electron density 1.25e20 $m^{-3}$
B (T) G (g/s) $\Delta P$ (bar)
3.8 9.9 0.85
2.8 10.3 0.68
1.8 10.95 0.4
Table 2: We perform simulations in which we vary the magnetic field strength by 1 T, leaving all other simulation parameters the same as in point 20 in table 4. Point 20 is represented by the middle row. In particular, these simulations use a 1 mm pellet with background plasma at approximately electron temperature 1220 eV and electron density 1.25e20 $m^{-3}$
@@ 4.5.1 Plasma gradients
Using PELOTON’s plasma cooling module, we computed plasma states after the injection of first 30 fragments (see Figure 9(a)), about 10 percent of a typical ITER injection, and used these states to estimate how plasma gradients affect the rocket acceleration. As we observe, after ablated material is deposited the plasma electron temperature remains practically constant, approximately 600 eV, within normalized flux coordinates from 0.9 to 0.5; plasma density also does not experience large gradients. As a result, plasma gradients changes the rocket acceleration by negligibly small amounts in this domain. Therefore, we estimate the upper bound of the plasma gradient effect by considering largest plasma gradients at the edge of the ablated material deposition with the flux coordinate of about 0.5.
(a) (a) Plasma electron density, $m^{-3}$ and (b) temperature (eV) after the injection of 30 SPI fragments.
@@ 5 Summary and future work
Appendix A Simulation Points Used for Ablation Rate and Rocket Acceleration in JET
Simulation No. Fragment Radius (mm) ne ($m^{-3}$) Te (eV) G (g/s) $\Delta P$ (bar) Acceleration ($ms^{-2}$)
1 0.625 5.7e19 554.67 1.135 0.16 9.7e4
2 0.48 6.1e19 786.065 1.190 0.23 1.8e5
3 0.31 6.1e19 920.823 1.198 0.26 3.1e5
4 0.19 6.1e19 997.636 .74 0.22 4.3e5
5 1.5 8.5e19 1530 23.1 0.68 1.7e5
6 1.0 8.2e19 1800 19.0 0.70 2.6e5
7 0.5 7.9e19 1982 8.9 0.48 3.6e5
8 0.25 7.8e19 2050 3.85 1.41 2.1e6
9 1.2 8.5e19 1230 11.1 0.52 1.6e5
10 0.8 8.5e19 1400 9.1 0.62 2.9e5
11 0.4 8.5e19 1550 4.5 0.53 4.9e5
12 0.2 8.5e19 1570 1.89 0.64 1.2e6
13 1.79 8.75e19 860.8 9.2 0.39 8.0e4
14 1.47 8.7e19 1175.5 13.8 0.55 1.4e5
15 0.4 8.5e19 1630.5 5.05 0.55 5.1e5
16 0.23 8.64e19 1602.5 2.45 0.80 1.3e6
17 1.0 1.8e20 1000 4.04 0.40 1.5e5
18 1.0 3.04e20 465.16 1.94 0.22 8.2e4
19 1.0 2.5e20 686.38 3.97 0.43 1.6e5
20 1.0 1.26e20 1219.33 10.3 0.68 2.5e5
Table 4: This table lists the simulation data points and results used to create function fits used in section 4.2.
Acknowledgment. This work has been supported in part by the ITER Organization and the ITER Disruption Mitigation Task Force via contract. The views and opinions expressed herein do not necessarily reflect those of the ITER Organization. The Swiss contribution was supported in part by the Swiss National Science Foundation.